OpenStudy (anonymous):

The speed of train A is 6 mph slower than the speed of train B. Train A travels 230 miles in the same time it takes train B to travel 360 miles. Find the speed of each train. I got B= 52 is this correct?

OpenStudy (anonymous):

Let train A travel at x-6 mph and train B travel at x mph for t hours and you get t(x-6)=230 miles and tx = 360 mile. Then x = 360/t and (x-6)=230/t. Substituting you get 360/t - 6 = 230/t -> 360 - 6t = 230 -> 6t = 130 -> t = 130/6 and x = 360*6/130 -> x = 216/13 mph